OpenStudy (anonymous):

x^-2+x^-1+1over x^-2-x eliminate the complex fractioons

OpenStudy (anonymous):

\[\frac{x^{-2}+x^{-1}+1}{x^{-2}-x}\] like that?

OpenStudy (anonymous):

multiply top and bottom by \(x^2\) to get \[\frac{x^{-2}+x^{-1}+1}{x^{-2}-x}\times \frac{x^2}{x^2}\] \[=\frac{1+x+x^2}{1-x^3}\]

OpenStudy (anonymous):

we are not done though, denominator factors \[\frac{1+x+x^2}{(1-x)(1+x+x^2)}\] and now you can cancel

OpenStudy (anonymous):

giving a final answer of \[\frac{1}{1-x}\]

OpenStudy (anonymous):

Yea exactly thank you for the help

OpenStudy (anonymous):

yw